# Multimodal LLMs Struggle with Basic Visual Network Analysis: a VNA Benchmark

### v1

@@ 5.3 Component Task Results
The top performing model across all evaluation metrics for both the component and isolate counting tasks was GPT-4 with no definitions included in the prompt. However, we note that the inclusion of definitions resulted in a substantial improvement of LLaVa’s MSE, which decreased from over 20,000 to below 11. GPT4 performed relatively well on the isolate counting task, and provided the correct number of isolates 67 of 100 times. We provide results in Table 3.
@@ 6 Discussion and Limitations
A limitation of this study is that there are a near large number of ways to visualize any given graph, and there are a myriad of parameters that are chosen which may impact LVM evaluation. It’s possible that something as trivial as changing node color could impact LVM evaluation. While the tasks we designed are representative of all possible graphs, they do evaluate mastery and understanding of the underlying network science concepts. The ways that graph visualization parameter selection and prompt engineering impact LVM performance on VNA tasks are clear and important avenues for future research.
@@ 7 Conclusion
We propose the task of zero-shot Visual Network Analysis to evaluate the performance of LVMs on graph analytics tasks. We create a benchmark that includes 5 tasks related to 3 core network science concepts: maximum degree, structural balance, and identifying components. We find, across all tasks LLMs struggled to identify and count the appropriate element of graphs—an essential skill in analyzing network data. We publicly release all data and generating code [^2] Repo will be made public shortly at: https://github.com/EvanUp/VNA_Benchmark.

### v2

@@ 5.3 Component Task Results
The top performing model across all evaluation metrics for both the component and isolate counting tasks was GPT-4 with no definitions included in the prompt. However, we note that the inclusion of definitions resulted in a substantial improvement of LLaVa’s MSE, which decreased from over 20,000 to below 11. gpt-4 performed relatively well on the isolate counting task, and provided the correct number of isolates 67 of 100 times. We provide results in Table 3.
@@ 6 Discussion and Limitations
A limitation of this study is that there are a large number of ways to visualize any given graph, and there are myriad parameters that are chosen which may impact LVM evaluation. It’s possible that something as trivial as changing node color could impact LVM evaluation. While the tasks we designed are not representative of all possible graphs, they do evaluate mastery and understanding of the underlying network science concepts. The ways that graph visualization parameter selection and prompt engineering impact LVM performance on VNA tasks are clear and important avenues for future research.
@@ 7 Conclusion
We propose the task of zero-shot Visual Network Analysis to evaluate the performance of LVMs on graph analytics tasks. We create a benchmark that includes 5 tasks related to 3 core network science concepts: maximum degree, structural balance, and identifying components. We find that across all tasks, LLMs struggled to identify and count the appropriate element of graphs—an essential skill in analyzing network data. We publicly release all generated data and ground-truth labels.